A table with smooth horizontal surface is turning at an angular speed Omega about its Axis. A groove is made on the surface along a radius of and a particle is gently placed inside the groove at a distance a from the centre. Find the speed of the particle as its distance from the centre becomes L.
Plz solve it from inertial frame or without using centrifugal force
1
Expert's answer
2020-12-28T09:00:31-0500

Answer

Now suppose v is the velocity of the particle with respect to the table at the instant the particle is at a distance x.When the particle is at a distance a from the centre, at that instant its velocity is zero.

Diagram can be shown as



Therefore acceleration

"a_r=x\omega^2=v\frac{dv}{dx}"

To taking this integration

"\int^v_0 vdv=\omega^2\int_a^Lxdx"

"v^2/2=\frac{\omega^2[x^2]_a^L}{2}"

So speed of particle is

"v=\omega\sqrt{L^2-a^2}"